# Feedback Connections in Quantum Reservoir Computing with Mid-Circuit Measurements

### v3

@@ I Introduction
In this paper, we introduce a novel feedback-driven quantum reservoir computing approach with mid-circuit measurements and classical feedforward operations during the coherence time of a qubit. In the restart based feedback protocol, as shown in Fig. 1 (c), each cycle requires $n$ shots. Thus, for an input sequence of training length $t$ tested over $k$ different reservoir realizations, the total number of shots scales with $t\times k\times n$. As $t$, $k$, or preferably $n$ grow, the cumulative cost of running these circuits becomes large. This motivates our exploration of feedback circuits operating within the coherence time of a qubit using novel classical feedforward methods. To isolate and rigorously quantify the effect of mid-circuit feedback, we perform a qubit reset after each time step. This ensures that no extraneous information (other than the intended feedback) propagates forward, allowing us to study the feedback coupling itself in a controlled manner. One can omit these resets to explore more complex scenarios in which correlations persist over time, but here our primary goal is to study feedback connections in isolation and thereby clarify their implications. For implementation, IBM superconducting systems have in particular demonstrated classical real-time feedforward operations [17], providing an opportunity to reduce the overall shot overhead for certain problems. We use the measurement strings obtained from a mid-circuit measurement as feedback for the next cycle. We can tune the feedback strength and the input strength through hyperparameters. We evaluate the memory performance by studying the short-term memory task and the predictive capabilities by predicting the Mackey-Glass system [18] and the expectation value of a spin in a chaotic Ising chain.
@@
Figure 1: QRC protocols: Each architecture is designed to process an input sequence ${s_{k}}$. The systems are initialized in the state $\rho_{0}$. A unitary operator $U(\cdot)$, constructed from parameterized Pauli rotations, governs the state evolution. Each $U(\cdot)$ represents the processing of one time step (cycle). (a) Graphical representation of a conventional restart-based QRC scheme. For each timestep, the architecture must be restarted and run from the beginning. This results in a quadratic runtime. (b) A rewind protocol that does not restart from the beginning, but only processes the last $\tau$ time steps. (c) A feedback controlled QRC protocol. Only one timestep is processed at a time, but the expectation value from the previous trial is stored in a buffer. (d) A mid-circuit measurement architecture that continuously monitors the quantum state. No restart is required. This architecture can be combined with weak measurements. (e) A feedback-driven QRC architecture with mid-circuit measurements. Feedback is incorporated by feeding the measurement strings $\textbf{m}_{k}$ back into the reservoir, a process that can be done at runtime.
@@ II-A Architecture
Here $\theta$ represents an angle parameter that determines the rotation applied in the two-qubit gate. Specifically, $\theta$ can take values from the set $\{a_{in}s_{k},a_{fb}m_{k-1}^{j}\}$, where $a_{in}$ and $a_{fb}$ are scaling parameters that control the input and feedback strength respectively. The structure of using $R$ gates along with the scaling parameters is commonly used in QRC models due a its compromise between being hardware efficient and generating substantial entanglement [9, 15, 19]. The full unitary $U(s_{k},\textbf{m}_{k-1})$ consists of an input $R$ gate, $N$ feedback $R$ gates, and a Haar-random unitary. The idea is illustrated for the case of $2$ qubits in Fig. 2. For the first time step $s_{0}$, a random vector $\textbf{m}_{-1}$ uniformly distributed from $\{-1,1\}^{N}$ is selected for time step 00. After unitary evolution, the system is measured projectively in the computational basis. We note that further research can also employ weak measurements, as seen in other QRC research [10, 15]. We reset our system to $\rho_{0}^{N}$ and feed the measurement string forward into the next unitary.
@@ IV-B QPU results
We now present experiments with our architecture on IBM hardware, specifically utilizing the dynamic circuit capabilities of ibm_marrakesh. Due to hardware constraints on the number of resets and conditional operations, we reduced the washout, training, and testing lengths to 5, 25, and 20 respectively. Additionally, time and hardware limitations made it infeasible to average over 128 Haar-random unitaries, so we selected a single fixed unitary. Below, we briefly describe how we generate a Haar-random unitary for the IBM device. A more detailed description of this algorithm is found in [33] and [34]:
@@ IV-C Echo state property analysis
The results of the ESP analysis are shown in Fig. 4. We see that all models aside from the proposed model with post-measurement resets, clearly exhibit the echo-state property. And even for the proposed model with post-measurement resets, we see that effect of the initial state becomes less prominent over time, although it does not completely vanish. This shows that all the QRC models investigated exhibit ESPs to at least some extent. To make sure that we did not stumble upon a special case, we found qualitatively similar results when we repeating the analysis with several different initializations of the fixed random parameters (e.g., $W_{in}$, $U_{\text{Haar}}$, etc.) and over different dimensions of the internal state of the reservoir.
@@ V Conclusion
In this paper, we propose a quantum reservoir computing scheme that uses mid-circuit measurements and feedback within the runtime of a qubit. Our initial experiments aim to determine whether feedback connection that happen during the coherence time of the qubit actually improves memory capacity, rather than merely introducing uncertainties due to measurement noise. We show that such an architecture with built-in resets has non-zero memory capacity for past inputs, indicating that feedback connections have an impact. In particular, peaks in memory capacity at specific feedback strengths are consistent with improved performance on prediction tasks involving both classical and quantum chaotic systems. Our ESP results show that most investigated QRC regimes exhibit the ESP, while our proposed model with post-measurement resets seems to have a mild form of the ESP. Further investigations are warranted to understand the importance of the ESP on the performance of QRC models.
@@ Appendix A Noise Test
We present preliminary noise tests for larger runs, accounting for the significant reductions in QPU runs. We use the same run statistics as in our numerical experiments. In our shot-based simulation, we introduce a depolarization channel after each unitary of the form:
@@
Figure 6: Noise Test: We show the noise effect of on our QRC regime on the NMSE of the Ising chain prediction tasks.

### v4

@@ I Introduction
In this paper, we introduce a novel feedback-driven quantum reservoir computing approach with mid-circuit measurements and classical feedforward operations during the coherence time of a qubit. In the restart-based feedback protocol, as shown in Fig. 1 (c), each cycle requires $n$ shots. Thus, for an input sequence of training length $t$ tested over $k$ different reservoir realizations, the total number of shots scales with $t\times k\times n$. As $t$, $k$, or preferably $n$ grow, the cumulative cost of running these circuits becomes large. This motivates our exploration of feedback circuits operating within the coherence time of a qubit using novel classical feedforward methods. To isolate and rigorously quantify the effect of mid-circuit feedback, we perform a qubit reset after each time step. This ensures that no extraneous information (other than the intended feedback) propagates forward, allowing us to study the feedback coupling itself in a controlled manner. One can omit these resets to explore more complex scenarios in which correlations persist over time, but here our primary goal is to study feedback connections in isolation and thereby clarify their implications. For implementation, IBM superconducting systems have in particular demonstrated classical real-time feedforward operations [17], providing an opportunity to reduce the overall shot overhead for certain problems. We use the measurement strings obtained from a mid-circuit measurement as feedback for the next cycle. We can tune the feedback strength and the input strength through hyperparameters. We evaluate the memory performance by studying the short-term memory task and the predictive capabilities by predicting the Mackey-Glass system [18] and the expectation value of a spin in a chaotic Ising chain.
@@
Figure 1: QRC protocols: Each architecture is designed to process an input sequence ${s_{k}}$. The systems are initialized in the state $\rho_{0}$. A unitary operator $U(\cdot)$, constructed from parameterized Pauli rotations, governs the state evolution. Each $U(\cdot)$ represents the processing of one time step (cycle). (a) Graphical representation of a conventional restart-based QRC scheme. For each timestep, the architecture must be restarted and run from the beginning. This results in a quadratic runtime. (b) A rewind protocol that does not restart from the beginning, but only processes the last $\tau$ time steps. (c) A feedback-controlled QRC protocol. Only one timestep is processed at a time, but the expectation value from the previous trial is stored in a buffer. (d) A mid-circuit measurement architecture that continuously monitors the quantum state. No restart is required. This architecture can be combined with weak measurements. (e) A feedback-driven QRC architecture with mid-circuit measurements. Feedback is incorporated by feeding the measurement strings $\textbf{m}_{k}$ back into the reservoir, a process that can be done at runtime. The expectation values $\langle Z\rangle_{k}$ are then computed in the post-processing step after all the circuits have been executed.
@@ II-A Architecture
Here $\theta$ represents an angle parameter that determines the rotation applied in the two-qubit gate. Specifically, $\theta$ can take values from the set $\{a_{in}s_{k},a_{fb}m_{k-1}^{j}\}$, where $a_{in}$ and $a_{fb}$ are scaling parameters that control the input and feedback strength respectively. The structure of using $R$ gates along with the scaling parameters is commonly used in QRC models due to its compromise between being hardware efficient and generating substantial entanglement [9, 15, 19]. The full unitary $U(s_{k},\textbf{m}_{k-1})$ consists of an input $R$ gate, $N$ feedback $R$ gates, and a Haar-random unitary. The idea is illustrated for the case of $2$ qubits in Fig. 2. For the first time step $s_{0}$, a random vector $\textbf{m}_{-1}$ uniformly distributed from $\{-1,1\}^{N}$ is selected for time step 00. After unitary evolution, the system is measured projectively in the computational basis. We note that further research can also employ weak measurements, as seen in other QRC research [10, 15]. We reset our system to $\rho_{0}^{N}$ and feed the measurement string forward into the next unitary.
@@ IV-B QPU results
We now present experiments with our architecture on IBM hardware, specifically utilizing the dynamic circuit capabilities of ibm_marrakesh. Due to hardware constraints on the number of resets and conditional operations, we reduced the washout, training, and testing lengths to 5, 25, and 20, respectively. Additionally, time and hardware limitations made it infeasible to average over 128 Haar-random unitaries, so we selected a single fixed unitary. Below, we briefly describe how we generate a Haar-random unitary for the IBM device. A more detailed description of this algorithm is found in [33] and [34]:
@@ IV-C Echo state property analysis
The results of the ESP analysis are shown in Fig. 4. We see that all models aside from the proposed model with post-measurement resets, clearly exhibit the echo-state property. And even for the proposed model with post-measurement resets, we see that effect of the initial state becomes less prominent over time, although it does not completely vanish. This shows that all the QRC models investigated exhibit ESPs to at least some extent. To make sure that we did not stumble upon a special case, we found qualitatively similar results when we repeating the analysis with several different initializations of the fixed random parameters (e.g., $W_{in}$, $U_{\text{Haar}}$, etc.) and over different dimensions of the internal state of the reservoir. During the course of our experiments there was no post-selection of the results, and all results from the different circuits composed of the various uniquely randomly generated $U_{Haar}$ matrices are included in the statistics .
@@ V Conclusion
In this paper, we propose a quantum reservoir computing scheme that uses mid-circuit measurements and feedback within the runtime of a qubit. Our initial experiments aim to determine whether feedback connection that happen during the coherence time of the qubit actually improves memory capacity, rather than merely introducing uncertainties due to measurement noise. We show that such an architecture with built-in resets has non-zero memory capacity for past inputs, indicating that feedback connections have an impact. In particular, peaks in memory capacity at specific feedback strengths are consistent with improved performance on prediction tasks involving both classical and quantum chaotic systems. Our ESP results show that most investigated QRC regimes exhibit the ESP, while our proposed model with post-measurement resets seems to have a mild form of the ESP.
@@ Appendix A Noise Test
We present preliminary noise tests for larger runs, accounting for the significant reductions in QPU runs. We use the same run statistics as in our numerical experiments. In our shot-based simulation, we introduce a depolarization (DP) channel after each unitary of the form:
@@
Figure 6: Noise Test: We analyze how noise in our QRC regime affects the NMSE in predicting the dynamics of the Ising chain.